OpenStudy (anonymous):

PLEASE HELP! Determine the equation of the line that passes through B(3,-1) and is perpendicular to the line defined by 2x+y-9=0.

OpenStudy (anonymous):

First you need to put the equation into y=mx + b form. Do you know how to do this?

OpenStudy (anonymous):

Are we suppose to take out the 0?

OpenStudy (anonymous):

Yes you just need to isolate y.

OpenStudy (anonymous):

y=2x - 9?

OpenStudy (anonymous):

Close your signs are wrong though

OpenStudy (anonymous):

which ones are wrong?

OpenStudy (anonymous):

Remember when you want to move stuff around you need to do the operation so you are taking away 2x from each side and adding 9 to each side so it cancels on the left and is moved to the right understand?

OpenStudy (anonymous):

It should be y = -2x + 9 know what I mean?

OpenStudy (anonymous):

why are we puting 2x as a negative though?

OpenStudy (anonymous):

Do you know how to do algebra?

OpenStudy (anonymous):

Yes but, im so confused right now

OpenStudy (anonymous):

Okay if you have y-5= 0 what does y= ?

OpenStudy (anonymous):

Oh shoot sorry! I get it now

OpenStudy (anonymous):

Since i want to transfer 2x and -9 to the other side i have to put it in the other operations as -2x and +9 right?

OpenStudy (anonymous):

=) that's okay. So now we have it in the form we need that was step one next we need the negative inverse of the slop because a line that is perpendicular will have a negative inverse slope do you know how to find this? Yes that is correct =).

OpenStudy (anonymous):

Would the -2x become as a 1/2? I dont know i did long time ago

OpenStudy (anonymous):

Correct so now you know the slope of your perpendicular line do you want to write out what it looks like so far?

OpenStudy (anonymous):

y= 1/2 + 9

OpenStudy (anonymous):

Wait the slope of are new equation is (1/2) in y = mx +b the slope is "m" and we don't know what b is yet.

OpenStudy (anonymous):

oh sorry!!!!! y=1/2x + 9

OpenStudy (anonymous):

close

OpenStudy (anonymous):

we don't know what b= so what should be in the place of b?

OpenStudy (anonymous):

do we plug in (3,-1) to variables of x and y?

OpenStudy (anonymous):

yes and solve for b because right now we have y = (1/2)x + b and we have a point on the line so we can solve for b like you said do you want to try that now?

OpenStudy (anonymous):

y= 1/2x + 9 (-1) =1/2(3) + 9 (-1) =1.5 + 9 (-1) = 10.5 -11.5? i dont think this is right... :(

OpenStudy (anonymous):

You don't know what b is you are trying to solve for b do you know what I mean?

OpenStudy (anonymous):

you have y = (1/2)x + b you need to plug in your points and solve for b.

OpenStudy (anonymous):

sorry! y=(1/2)x + b (-1) = (1/2)(3) + b (-1) = (1.5) + b (-1) - (1.5) = b -2.5 = b

OpenStudy (anonymous):

Looks good to me so what is your equation?

OpenStudy (anonymous):

y= 1/2x - 2.5?

OpenStudy (anonymous):

Correct =) good job

OpenStudy (anonymous):

but it says in my book that it's x-2y-5=0....?

OpenStudy (anonymous):

multiply everything by 2 and solve for zero to get it into that form. Doing this isn't changing the equation all the points will be the same.

OpenStudy (anonymous):

That should say + 5

OpenStudy (anonymous):

So the equation can be what we just did right? it doesn't have to be the form in which my book did it?

OpenStudy (anonymous):

Not unless the teacher wants you to put it into that form.

OpenStudy (anonymous):

Are you sure the one in the book doesn't say x - 2y + 5 = 0 ?

OpenStudy (anonymous):

how do i put it in that form though? and no it doesnt say +5 :(

OpenStudy (anonymous):

oh my bad sorry I got confused it should say -5 haha

OpenStudy (anonymous):

Anyway you just multiply everything in the equation had by 2 and then you move everything to the left side and set the right to zero know what I mean?

OpenStudy (anonymous):

not exactly...

OpenStudy (anonymous):

2*y = 2*((1/2)x + 2.5) what does it look like after doing this?

OpenStudy (anonymous):

2y=y +5 and the you put the x and 5 on the other side?

OpenStudy (anonymous):

right =)

OpenStudy (anonymous):

so its a 2y= -x -5?

OpenStudy (anonymous):

wait look at what you typed

OpenStudy (anonymous):

you need 2y - x - 5 = 0

OpenStudy (anonymous):

So you don't put the equals sign?

OpenStudy (anonymous):

what did the equation in the book say?

OpenStudy (anonymous):

x-2y-5=0

OpenStudy (anonymous):

okay so if you have 2y = x - 5 then you need to move the stuff on the right onto the left, and then you would have is -x+ 2y + 5 = 0 but what you need is x - 2y - 5= 0 so you multiply the equation by -1 and then you are done. Honestly these steps are pointless but it is just the way you make it look like the one in the book.

OpenStudy (anonymous):

This is really confusing but i think i get it.... Why did you multiply by 2 in the first place though from the previous one?

OpenStudy (anonymous):

In order to get the same coefficients

OpenStudy (anonymous):

Oh alright, THANK YOU SO MUCH! :)

OpenStudy (anonymous):

No problem you solved it =)